Find the minimum number of intervals required to evaluate int (e^(-x^2 +1))dx with an accuracy of 0.5 × 10^(-4) , by using Trapezoidal rule.
1
Expert's answer
2020-04-13T19:05:22-0400

To find the minimum number "n" of intervals, we need to solve the inequality


"E\leq\frac{(b-a)^3}{12n^2}[\text{max}|f"(x)|],a\\leq x\\leq b."

Therefore, first find the second derivative:


"\frac{d}{dx}\big[e^{-x^2+1}\big]=-2xe^{1-x^2}=f'(x),\\ \space\\
\frac{d}{dx}\big[-2xe^{1-x^2}\big]=2e^{1-x^2}(2x^2-1)=f"(x)."

Assume that the integration limits are "a=-1, b=2" because they are not present in the condition. Therefore:

"f"(-1)=2\\
f"(2)=0.7\\\\\nf"(0)=-5.43656...=\text{max}[f"(x)].\\\\"

Find "n":


"0.5\cdot10^{-4}\leq\frac{(2-(-1))^3}{12n^2}[5.27]=\frac{27}{12n^2}\cdot5.4366,
\\ \space\\
n^2\geq244647,\\n\geq494.62\approx496."